Lenox is planning to read 3 books in 5 weeks. The average number to read the same number books is 184. He wants to read the same number of pages everyday except on Friday, when he doesn't plan to read. How many pages should he read per day to finish the 3 books in 5 weeks?
Answer by checkley71(8403)

3*184=552 ARE THE TOTAL NUMBER OF PAGES TO BE READ.
5 WEEKS MINUS FRIDAYS (7-1)*5=6*5=30 DAYS OF READING
552/30=18.4 PAGES PER DAY WILL DO IT.